Use the substitution method to solve m - n = 7 and 3m + 2n = 6.
Answer by fractalier(6550)

Okay, from
m - n = 7 and 3m + 2n = 6,
we solve the first one for m,
m = n + 7 \and substitute that into the second one, like this...
3(n+7) + 2n = 6
3n + 21 + 2n = 6
5n + 21 = 6
5n = -15
n = -3
Now m = n+7 = -3 + 7 = 4